OpenStudy (anonymous):

MEDAL :D #help A used car has a value of $15,250 when it is purchased in 2012. The value of the car depreciates at a rate of 7.5% per year. Write an exponential function that models the value of the car, y, over x years. To the nearest cent, what will be the value of the car after eight years?

OpenStudy (anonymous):

y = 15250*(1 - 0.075)^x after 8 years we have x = 8 y = 15250*(1 - 0.075)^8 y = $8,173.42 after 8 years the value of the car will be $8,173.42.